# Visualizing, Rethinking, and Mining the Loss Landscape of Deep Neural Networks

### v1

@@ Appendix B Demo Code
If we aim to obtain the largest Hessian eigenvalue and corresponding eigenvector, the power iteration method could be utilized. Specifically, given a random vector $v$, we could keep calculating the “Jacobian Vector Product” by $v\leftarrow\frac{Hv}{||Hv||}$. After convergence, we could obtain the largest eigenvalue and corresponding eigenvector. However, calculating the smallest eigenvalue and corresponding eigenvector is slightly complex. We utilize the package of sparse.linalg.eigsh to accomplish this goal. eigsh accepts as the input a square operator representing the operation $Hv$, where $H$ is real symmetric or complex Hermitian. This condition is satisfied by the Hessian matrix and the trick of “Jacobian Vector Product”. The demo code is listed in Code 1, which also utilizes the torch.autograd and LinearOperator. We omit the introduction of these functions and these could be found on the web easily. The function parameter “which” determines the types of eigenpairs. If “which” is set as “LA”/“SA”, the function calculates the largest/smallest algebraic eigenvalues and corresponding eigenvectors. We calculate the largest top $k=10$ eigenpairs and smallest $k$ eigenpairs and denote the eigenvectors as “P.E.x” and “N.E.x” respectively. The value of “x” ranges from 1 to 10, and a smaller “x” refers to that the corresponding eigenvalue has a larger absolute value.
@@
Listing 1: Hessian eigenvectors by sparse.linalg.eigsh
1import torch
2import torch.nn.functional as F
3from torch.autograd import grad
4from scipy.sparse.linalg import LinearOperator, eigsh
5
6inputs = None
7labels = None
8model = None
9n_params = None # number of total trainable parameters
10
11def hessian_vector_product(vector):
12 loss = F.cross_entropy(model(inputs), labels)
13 grad_params = grad(loss, model.parameters(), create_graph=True)
14 flat_grad = torch.cat([g.view(-1) for g in grad_params])
15 grad_vector_product = torch.sum(flat_grad * vector)
16 hvp = grad(grad_vector_product, model.parameters(), retain_graph=True)
17 return torch.cat([g.contiguous().view(-1) for g in hvp])
18
19def matvec(v):
20 v_tensor = torch.tensor(v, dtype=torch.float32)
21 return hessian_vector_product(v_tensor).cpu().detach().numpy()
22
23eigenvalues, eigenvectors = eigsh(
24 LinearOperator((n_params, n_params), matvec=matvec),
25 k=10, tol=1e-3, which="SA", return_eigenvectors=True
26)
@@ B.2 Calculating the eigenvalue density of the Hessian by PyHessian
PyHessian [68] is a package that calculates the statistics of the Hessian matrix for DNNs, which includes: (1) the most top-k largest eigenvalues and corresponding eigenvectors that utilize the power iteration method; (2) the trace of the Hessian matrix that could be estimated by $\mathbb{E}_{\epsilon}[\epsilon^{T}H\epsilon]$ where $\epsilon$ is element-wisely sampled from $\mathcal{N}(0,1)$; (3) the eigenvalue density of eigenvalues approximated by the algorithm of Stochastic Lanczos Quadrature (SLQ) [46]. We download the source codes of PyHessian and utilize the “PyHessian” class and the “density_generate” function to plot the eigenvalue density. The demo code is listed in Code 2.
Listing 2: Hessian density of eigenvalues by PyHessian
1import torch.nn.functional as F
2from pyhessian import PyHessian, density_generate
3from matplotlib import pyplot as plt
4
5model = None # model
6loader = None # DataLoader
7
8algo = PyHessian(
9 model=model,
10 criterion=F.cross_entropy,
11 data=None,
12 dataloader=loader
13)
14
15eigvs, weights = algo.density()
16xs, ys = density_generate(eigvs, weights)
17plt.semilogy(xs, ys + 1.0e-7)
@@
As shown in Eq. 1, the second-order approximation formulation is $\delta\mathcal{L}=\mathcal{L}(\theta+\lambda\epsilon)-\mathcal{L}(\theta)\approx\lambda\epsilon^{T}g_{\theta}+\frac{1}{2}\lambda^{2}\epsilon^{T}H_{\theta}\epsilon$. This could be viewed as a quadratic function of $\lambda$, i.e., $\frac{1}{2}a\lambda^{2}+b\lambda$ with $a=\epsilon^{T}H_{\theta}\epsilon$ and $b=\epsilon^{T}g_{\theta}$. Hence, given a model point $\theta$, we could calculate $a$ and $b$ by the trick of “Jacobian Vector Product” again. For each given $\lambda$, we could sample multiple groups of $\epsilon$ and then plot the distribution of $\delta\mathcal{L}$. The demo code is listed in Code 3.
@@
Listing 3: Simulating the second approximation as in Eq. 1
1import torch
2import torch.nn.functional as F
3
4inputs = None
5labels = None
6model = None
7
8def get_params_grad(model):
9 params = []
10 grads = []
11 for _, param in model.named_parameters():
12 if not param.requires_grad:
13 continue
14 params.append(param)
15 grads.append(0. if param.grad is None else param.grad + 0.)
16 return params, grads
17
18def hessian_vector_product(gradsH, params, v):
19 hv = torch.autograd.grad(
20 gradsH, params, grad_outputs=v, only_inputs=True, retain_graph=True)
21 return hv
22
23def calculate_derivatives(epsilon_model):
24 model.zero_grad()
25 loss = F.cross_entropy(model(inputs), labels)
26 loss.backward()
27 params, grads = get_params_grad(model)
28 epsilon_vec = [p.detach() for p in epsilon_model.parameters()]
29 fd, sd = 0.0, 0.0
30 fd = sum([(p.data * g.data).sum() for p, g in zip(epsilon_vec, grads)]).item()
31 hv = hessian_vector_product(grads, params, v=epsilon_vec)
32 sd = sum([(p.data * g.data).sum() for p, g in zip(epsilon_vec, hv)]).item()
33 model.zero_grad()
34 return fd, sd
35
36
37lamb = 0.0
38vs = []
39for _ in range(100):
40 epsilon_model = None # sample Gaussian noise model
41 fd, sd = calculate_derivatives(epsilon_model)
42 vs.append(lamb * fd + 0.5 * lamb * lamb * sd)

### v2

@@ Abstract
keywords:
Deep neural networks , Loss landscape visualization , Monotonic Linear Interpolation
journal: Knowledge-Based Systems
[…]
@@ Appendix B Demo Code
If we aim to obtain the largest Hessian eigenvalue and corresponding eigenvector, the power iteration method could be utilized. Specifically, given a random vector $v$, we could keep calculating the “Jacobian Vector Product” by $v\leftarrow\frac{Hv}{||Hv||}$. After convergence, we could obtain the largest eigenvalue and corresponding eigenvector. However, calculating the smallest eigenvalue and corresponding eigenvector is slightly complex. We utilize the package of sparse.linalg.eigsh to accomplish this goal. eigsh accepts as the input a square operator representing the operation $Hv$, where $H$ is real symmetric or complex Hermitian. This condition is satisfied by the Hessian matrix and the trick of “Jacobian Vector Product”. The demo code is listed in Code, which also utilizes the torch.autograd and LinearOperator. We omit the introduction of these functions and these could be found on the web easily. The function parameter “which” determines the types of eigenpairs. If “which” is set as “LA”/“SA”, the function calculates the largest/smallest algebraic eigenvalues and corresponding eigenvectors. We calculate the largest top $k=10$ eigenpairs and smallest $k$ eigenpairs and denote the eigenvectors as “P.E.x” and “N.E.x” respectively. The value of “x” ranges from 1 to 10, and a smaller “x” refers to that the corresponding eigenvalue has a larger absolute value.
@@ B.2 Calculating the eigenvalue density of the Hessian by PyHessian
PyHessian [24] is a package that calculates the statistics of the Hessian matrix for DNNs, which includes: (1) the most top-k largest eigenvalues and corresponding eigenvectors that utilize the power iteration method; (2) the trace of the Hessian matrix that could be estimated by $\mathbb{E}_{\epsilon}[\epsilon^{T}H\epsilon]$ where $\epsilon$ is element-wisely sampled from $\mathcal{N}(0,1)$; (3) the eigenvalue density of eigenvalues approximated by the algorithm of Stochastic Lanczos Quadrature (SLQ) [70]. We download the source codes of PyHessian and utilize the “PyHessian” class and the “density_generate” function to plot the eigenvalue density.
@@
As shown in Eq. 1, the second-order approximation formulation is $\delta\mathcal{L}=\mathcal{L}(\theta+\lambda\epsilon)-\mathcal{L}(\theta)\approx\lambda\epsilon^{T}g_{\theta}+\frac{1}{2}\lambda^{2}\epsilon^{T}H_{\theta}\epsilon$. This could be viewed as a quadratic function of $\lambda$, i.e., $\frac{1}{2}a\lambda^{2}+b\lambda$ with $a=\epsilon^{T}H_{\theta}\epsilon$ and $b=\epsilon^{T}g_{\theta}$. Hence, given a model point $\theta$, we could calculate $a$ and $b$ by the trick of “Jacobian Vector Product” again. For each given $\lambda$, we could sample multiple groups of $\epsilon$ and then plot the distribution of $\delta\mathcal{L}$.